OpenStudy (anonymous):

HELP! MEDAL :D Solve. Express the answer as an inequality statement or the empty set. x - 3 ≤ 4x and 2x + 6 < 18 @Directrix @SolomonZelman @myininaya @mathstudent55 @texaschic101 @iambatman @whpalmer4 @sammixboo @freckles

OpenStudy (freckles):

can you solve \[x-3 \le 4x \]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x ≤ 1x?

OpenStudy (freckles):

We are going to solve both inequalities and take the intersection of the solutions for the inequalities

OpenStudy (freckles):

hmm... well that isn't solved for x

OpenStudy (freckles):

we have x-3<=4x I would first subtract x on both sides -3<=3x

OpenStudy (freckles):

do you know what to do here?

OpenStudy (freckles):

you have \[-3 \le 3x\]

OpenStudy (freckles):

to get x by itself you are going to do what to both sides?

OpenStudy (anonymous):

divide

OpenStudy (freckles):

by 3 right?

OpenStudy (freckles):

which will give?

OpenStudy (anonymous):

wait why and who did you subtract the x at the beginning

OpenStudy (anonymous):

how*

OpenStudy (freckles):

] that is how you undo addition you subtract...